OpenStudy (anonymous):

Describe how to transform (^6sqrtx^5)^7 into an expression with a rational exponent.

OpenStudy (anonymous):

\[\LARGE (\sqrt[6]{x^5{}})^{7}\]

OpenStudy (jdoe0001):

\(\huge a^{\frac{n}{m}} = \sqrt[m]{a^n}\)

OpenStudy (anonymous):

What?

OpenStudy (jdoe0001):

\(\Large \bf \left(\sqrt[6]{x^5}\right)^7\implies \sqrt[6]{(x^5)^7}\implies \sqrt[6]{x^{35}} \implies x^{\frac{35}{6}}\)

OpenStudy (anonymous):

Oh okay thank you :)